Question No. 01

 

 

PakMonoG’s inverse demand function is P = 100 – 2Q and cost function is TC = 10 + 2Q,

Where Q is quantity in units and P price in PKR.

a.      Determine the profit-maximizing price, quantity and profit (or loss) of PakMonoG.  

b.     Given your calculations in (a), illustrate the demand, marginal revenue and marginal cost curves of the firm in a graph.  

c.      If we were to compare PakMonoG with a perfect competitive firm in the market, are there differences in characteristics of the two structures?  

What are welfare implications? Is total societal welfare of the firm higher or lower than that of a competitive firm? Support your answer using the graph in (b) above


1
Expert's answer
2021-05-30T14:19:23-0400
"MC=\frac{\delta TC}{\delta Q}=2"

"MC=p"

"Q=49"

"\pi=TR-TC=98-108=-10"

"\pi=2Q-10-2Q=-10"

"TR=-2Q^2+98Q-10"

"MR=-4Q+98"